OpenStudy (anonymous):

Simplify the following rational function: 2x-1/1-2x

jimthompson5910 (jim_thompson5910):

\[\large \frac{2x-1}{1-2x}\] \[\large \frac{2x-1}{-1(-1+2x)}\] \[\large \frac{2x-1}{-1(2x-1)}\] \[\large \frac{\cancel{2x-1}}{-1\cancel{(2x-1)}}\] \[\large \frac{1}{-1}\] \[\large -1\] So \[\large \frac{2x-1}{1-2x}=-1\]

OpenStudy (anonymous):

thanks